ON MEASUREABLE SETS OF POSITIVE MEASURE by Mukul Pal and Mrityunjoy Nath, Kalyani, India

Abstract. In this paper, along with other results, we extend a result of M.Pal. Mathematics subject classification: 28A10 Keywords: Euclidean space, Lebesgue measure